How many gallons of a 2% salt solution must be mixed with 15 gallons of a 8% solution to obtain a 5% solution?
Answer by richwmiller(17219)

0.02*x+0.08*15=0.05(15+x)
0.02*x+1.2=0.75+0.05x
0.02x-0.05x=0.75-1.2
-0.03x=-0.45
x=15 gal of 2% is added to 15 gal of 8% to get 30 gal at 5%
check
0.02*15+0.08*15=0.05(15+15)
0.3+1.2=0.05(30)
1.5=1.5
ok
Note that 5 is the average of 2 and 8 so we need equal amount of 2% and 8%